OpenStudy (anonymous):

Write the equation in slope-intercept form. What are the slope and y-intercept? -11x + 9y = -12

OpenStudy (anonymous):

Do you know what the slope-intercept form is?

OpenStudy (anonymous):

A. y = 11/9x - 4/3 slope: 11/9 y-intercept: 4/3 B. y = -11/9x + 4/3 slope: 11/9 y-intercept: 4/3 C. y = 11/9x + 4/3 slope: 4/3 y-intercept: 11/9 D. y = 11/9x - 4/3 slope: 11/9 y-intercept: -4/3

OpenStudy (anonymous):

So, those are your choices. Do you know the slope-intercept form?

OpenStudy (anonymous):

y = 11/9x -4/3

OpenStudy (anonymous):

Oh come on. Don't just hand out answers like candy. That doesn't help anyone.